The sum of the interior angles of a polygon is 8y. If Y is 15 greater than the number of sides of the polygon, how many sides does the polygon have?
Answer by stanbon(57377)

The sum of the interior angles of a polygon is 8y. If Y is 15 greater than the number of sides of the polygon, how many sides does the polygon have?
---
Let number of sides be "n":
sum of interior angles = (n-2)180 = 8Y
y = n+15
---
Substitute for "y" and solve for "n"
-----------------------
(n-2)180 = 8(n+15)
----
(n-2)180 = 8n+120
180n-360 = 8n+120
172n = 480
-----
That does not turn out to be a whole number.
Please check the wording of your problem.
I have some suspicion about that "8Y".
----
Cheers,
Stan H.
==============